OpenStudy (loser66):

Statistics need help, please If \(Y_1,Y_2,\cdots ,Y_n\)~N\((\mu, \sigma^2)\), then, \(\dfrac{\overline Y -\mu}{\dfrac{\sigma}{\sqrt n}}=Z\)~N\((0,1)\)

OpenStudy (loser66):

And then, on Sample mean distribution, I have \(Z=\dfrac{\overline Y -\mu}{\sigma}\)~N(0,1) My question is : where is \(\sqrt n\)?

OpenStudy (loser66):

We will use the second Z to make a chi-square distribution. But it is unclear to me. Please, explain

OpenStudy (loser66):

@Zarkon Please, explain me

OpenStudy (caozeyuan):

@Loser66 , so if you have a normal distribution with mean mu and variance sigma squared, you can transfer your distribution to standard normal but caculating Z score which is (X-mu) over sigma, where X is original distribution data

OpenStudy (caozeyuan):

If you have sampling distribution of sample mean, with know population variance sigma squared, then according to CLT, you have the (X-mu)/standard error is normally distributed with mean 0 and variance 1

OpenStudy (caozeyuan):

so basicallyZ score is only used to transfer a given normal distribution to standard normal and the formula with sigma oveer sqrt(n) is to be used if you have a sapling distribution of sample mean

OpenStudy (loser66):

Thanks for the explanation, but my question is about \(\sqrt n\) For both, they are almost the same but \(\sqrt n\), both are standard normal, right? but they are NOT equal, right? I meant \(Z=\dfrac{\overline Y-\mu}{\sigma /\sqrt n}\neq \color {red}{Z}=\dfrac{\overline Y-\mu}{\sigma}\) That is only one way made it clear to me. If they are equal, I ...... don't understand

OpenStudy (zarkon):

this \[\dfrac{\overline Y-\mu}{\sigma}\] is not standard normal (unless n=1) but \[\dfrac{Y_i-\mu}{\sigma}\] is standard normal

OpenStudy (loser66):

Yes, I see the different between them, the first one is \(\sqrt n(\overline Y-\mu)/\sigma \) and the second one is \((Y_i-\mu)/\sigma\) How do they relate to each other?

OpenStudy (zarkon):

only that they are both standard normal (provided your \(Y_i\) are independent)

OpenStudy (loser66):

But they are different, right?

OpenStudy (zarkon):

What do you mean by different? They are both random variables that have the same distribution

OpenStudy (zarkon):

do they take the same value for each realization...no, but that really doesn't matter

OpenStudy (loser66):

Like this \(Y_1, Y_2,\cdots ,Y_n\)~N(0,1) but they are different since they are independent even though all are standard normal distribution. Same as those Zs

OpenStudy (zarkon):

like I said they will have different realizations

OpenStudy (zarkon):

I'm not sure what you are trying to get at?

OpenStudy (loser66):

If they use \(Z_1=\sqrt n\dfrac{\overline Y -\mu}{\sigma}\) ~ N (0,1) and \(Z_2=\dfrac{Y_i -\mu}{\sigma}\)~ N (0,1) then I am ok. I understand they are 2 different standard normal distributions. But they use the same Z. It makes me think they are only 1 normal distribution but represented in 2 forms.

OpenStudy (zarkon):

they are different \[P(Z_1=Z_2)=0\]

OpenStudy (zarkon):

again, provided n>1

OpenStudy (loser66):

One more question :) I don't get the last result from the solution \(m_z (t) = e^t^2/2 from this

OpenStudy (zarkon):

about the mean and variance being 0 and 1 respectively?

OpenStudy (loser66):

\(m_{(Y-\mu)}(t/\sigma)\) = ?

OpenStudy (zarkon):

you don't understand why it is equal to \[e^{(t/\sigma)^2(\sigma^2/2)}\]?

OpenStudy (loser66):

yes, please

OpenStudy (zarkon):

you understand why \[m_{(Y-\mu)}(t)=e^{(t^2)(\sigma^2/2)}\]?

OpenStudy (loser66):

oh, just replace t by t/sigma, right?

OpenStudy (zarkon):

yes

OpenStudy (loser66):

ha!! :) Thanks a lot. Ah!! I am struggling with statistics.

OpenStudy (zarkon):

what book are you using?

OpenStudy (loser66):

Mathematical Statistics with Applications 7th Edition

OpenStudy (zarkon):

Mendenhall

OpenStudy (loser66):

Yes, Dennis D. Wackerly , William Mendenhall III and Richard L. Scheaffer

OpenStudy (zarkon):

I'm teaching out of that book this semester ;)

OpenStudy (zarkon):

we are on section 6.3...starting 6.4 tomorrow

OpenStudy (loser66):

Really? I stuck everywhere because I know nothing about statistics. My Prof will teach me 7.2 or 7.3 next week but I have to study the parts I don't get.

OpenStudy (zarkon):

ic

OpenStudy (loser66):

Thank you so much. :)

OpenStudy (zarkon):

no problem

OpenStudy (loser66):

I am sorry for bothering you again. I would like to know how to find cdf /pdf of chi-square on Ti 83. My Prof said something about the table, but I know there is a way we can find it out by calculator, just not know how to yet. :) Please, teach me.

OpenStudy (zarkon):

hit 2nd - vars the \(\chi^2\) are numbers 7 and 8

OpenStudy (loser66):

I use Ti 83 plus, so that I don't have upper /lower and df But I guess!! :) 2nd +Vars then (0, 5, 7) where 0 is lower, 5 is upper limit and 7 is degree of freedom, right?

OpenStudy (loser66):

\(Y_1\)~N (0,1) \(Y_2\)~N (0,1) \((Y_1-Y_2)\)~ N( 0,2) and then \(\dfrac{Y_1-Y_2}{\sqrt 2}\) ~ N (0,1) why?

OpenStudy (loser66):

I understand why Y1-Y2 ~ N (0,2), just the last one.

OpenStudy (zarkon):

if you have a random variable X and you divide by its standard deviation then you get a random variable with s.d. 1 2 is the variance and \(\sqrt{2}\) is the standard deivation

OpenStudy (loser66):

Perfect. Thanks a ton